# PECAN: Personalizing Robot Behaviors through a Learned Canonical Space

### v1

@@ 6.1. Learning User-Friendly Canonical Spaces
Experimental Setup. Participants in this study interacted with a $6$-DoF UR5 robot arm in three manipulation tasks: (i) handing over a plate to the user (Handover Plate), (ii) placing a cup in front of the user (Place Cup), and (iii) pouring coffee into the cup (Pour Coffee). In each task, the robot’s style was defined by the distance it maintained from the user. At one extreme, the robot could follow a straight-line path that comes very close to the user, while at the other extreme, it could take a curved path that stays as far from the user as possible. To personalize the robot’s trajectory, each user interacted with an interface containing a canonical space of the robot’s styles. The interface featured a task selection menu, a slider for choosing the latent style (as shown in Fig. 4-Left), and buttons to execute the trajectory corresponding to the chosen style in a Pybullet simulation and on the real robot. We pre-programmed the robot to pick up the objects for each task (i.e., plate, cup, or kettle) from their initial positions. The robot then performed the tasks according to the style selected by the user. See videos here: https://youtu.be/wRJpyr23PKI
@@
Participants performed the three tasks in the order given above. In the first two tasks, Handover Plate and Place Cup, they could preview the robot’s trajectory in the Pybullet simulation before executing it in the real world. However, they did not have this option in the third task of Pour Coffee and had to find their target style based on their experience of using the interface in the first two tasks. We did this to determine if users could build an accurate understanding of the canonical space, enabling them to effectively transfer it to new tasks. Therefore, we treated the first two tasks as familiarization tasks and evaluated the user performances in the Pour Coffee task. We anticipated that a consistent and monotonic canonical space would be easier for users to learn and apply across new tasks without the need to visualize the robot’s behavior.
@@
Independent Variables. We compared our proposed approach (PECAN) to a state-of-the-art baseline for generating latent style representations (SeGMA). To train these methods, we provided 3 demonstrations each in the Handover Plate and Place Cup tasks. Two of those demonstrations represented the extreme styles in that task — trajectories with the maximum and minimum distance to the user. Additionally, we asked each participant to provide $2$ demonstrations in the Pour Coffee task, one for each of the extreme styles. In total, we trained PECAN and SeGMA using a dataset of 8 demonstrations with labels for the trajectories with extreme styles. For PECAN, we assigned the same label to trajectories with similar styles across the tasks. On the other hand, we provided the same label for trajectories in the same task when training SeGMA.
@@
Table 1. Survey questions (Likert scales with $7$-option response format). We grouped questions into five scales and tested their reliability using Cronbach’s $\alpha$. The reliability scores presented in the table are based on the responses recorded in the robot user study.
@@
Participants and Procedure. We recruited $14$ participants ($3$ female, average age $28\pm 5$ years) from the Virginia Tech community. Participants gave informed written consent under IRB #23-1237. At the start of the experiment, users kinesthetically guided the robot arm to demonstrate trajectories for the extreme styles in the Pour Coffee task. We then trained both methods on the user demonstrations along with the six previously collected demonstrations. Users interacted with the robot in all three tasks — once with each method. They completed the tasks in a fixed order but the ordering of the methods was counterbalanced: half of the users started with PECAN and the other half started with SeGMA.
In the Handover Plate and Place Cup tasks, users personalized the robot’s trajectory once to match distinct target styles. Then, in the Pour Coffee task, users personalized the robot’s motion twice to achieve two additional target styles. We randomly sampled the four target styles for each user. For each style, users had $3$ attempts to perform the task on the real robot while ensuring that its trajectory stayed within a small tolerance of the desired distance. To help users gauge the actual style of the robot’s trajectory, we displayed its maximum distance from the user on the interface.
Dependent Variables. For each task we recorded the number of attempts that users took to achieve the target style on the real robot (Real Attempts). We also recorded the number of times users visualized the latent styles in simulation before executing them in the real world (Sim Attempts) for the Handover Plate and Place Cup tasks. A higher number of attempts indicated that it was difficult for users to identify their desired styles in the learned canonical space. Specifically, in the Pour Coffee task, where users did not have the option to simulate the styles, a higher number of attempts indicated that the interface was not intuitive and consistent across tasks. We also measured the error in the distance (Style Error) and final position (Task Error) of the trajectories executed by users in their last attempt for each target style.
@@
Figure 5. Objective and subjective results from the robot user study. (Left) User applies the same latent value $z_{\theta}$ from Place Cup to Pour Coffee expecting a similar style across both tasks. PECAN produces trajectories with similar distances $\theta$ for Place Cup and Pour Coffee, but SeGMA generates a straight line trajectory for Pour Coffee. (Top Right) When using PECAN participants had lower task error, style error, and fewer real attempts (t(13) = -3.12, p < 0.01) performing new tasks without visual information. (Bottom Right) Subjectively participants prefer (t(13) = 2.55, p < 0.05) working with PECAN compared to SeGMA, as they found PECAN more intuitive (t(13) = 3.35, p < 0.01), and easy (t(13) = 2.68, p < 0.05) especially without visual information (t(13) = 2.76, p < 0.05). An asterisk (*) denotes statistical significance.
@@
In the second user study we determined the pros and cons of directly selecting the robot’s style compared to indirect methods that estimate the style from user feedback (e.g., ranking robot trajectories). Participants were presented with two fundamentally different approaches for modifying an autonomous car’s driving style — our direct approach, PECAN, and an established approach for indirectly learning from human preferences (Sadigh et al., 2017).
Independent Variables. Specifically, we compared PECAN to an active preference-based learning approach, which we refer to as APReL. We implemented this approach based on the code provided in (Bıyık et al., 2022b). At each step, APReL presented two trajectories to the users, each representing a different style, and asked them to choose their preferred trajectory. Based on their choices, APReL inferred the individual user styles. It strategically selected these trajectories to maximize the information it gained about the user’s style from their choice. For example, showing trajectories with distinctly different speeds, such as one high-speed and one low-speed, is more informative than showing two trajectories with similar speeds. In our implementation, APReL selected the trajectories from a discrete space of $176$ uniformly sampled styles. We trained PECAN with $24$ demonstrations: $8$ demonstrations corresponded to the extreme styles and $16$ were sampled randomly, as explained in Section 5. It is important to note that, unlike our approach, APReL has direct access to the features that parameterize the driving style of the autonomous car. Therefore, we might expect this baseline to outperform our approach because it knows the actual styles, while our approach must learn these styles from user demonstrations.
@@
Dependent Variables. For the indirect approach (APReL) we recorded the total number of queries that users had to answer for personalizing the car’s behavior. For the direct approach (PECAN) we counted the total number of points that users had to visualize for finding their preferred style. We collected the subjective responses of users for the same scales as in Table 1. We also added another scale for measuring if users perceived that they required fewer attempts (clicks or queries) to personalize the robot as they gained more experience with each method (Learn).
@@
Figure 6. Objective and subjective results for the second user study. (Left) The average error in the style of the personalized car trajectory in both tasks. (Center) The total number of attempts (queries or clicks) required by users to personalize the car in both tasks. (Right) Users found both the direct and indirect approaches to be intuitive, accurate, and easy to use. In particular, they perceived that they needed fewer clicks to find their desired style as they gained more experience with our direct approach.
@@
Figure 7. Objective and subjective results for the follow-up study comparing direct and indirect approaches for personalizing robot styles with experienced users. (Left) The average error in the style of the personalized car trajectory in both tasks. (Center) The average number of attempts (queries or clicks) required by users to personalize the car for each target style across both tasks. After practicing with both approaches for $10$ minutes, users were able to achieve their target style more accurately with our direct approach, PECAN, and needed fewer attempts to do so than the indirect baseline, APReL. (Right) Subjectively, experienced users found PECAN to be more intuitive, accurate, and easy to use than APReL. Consequently, they preferred it over the indirect baseline for personalizing the car’s trajectory across tasks.
@@
Results. The results of the follow-up study are shown in Figure 7. After practicing with both approaches, we observed that users required significantly fewer clicks or queries to personalize the autonomous car’s style with PECAN than with APReL. A two-tailed paired t-test showed a statistically significant difference ($t(9)=-3.28$, $p<0.01$) between the average number of attempts required by users per style with PECAN ($M=2.1$, $SE=0.33$) and APReL ($M=3.9$, $SE=0.48$). Subjectively, experienced users found our direct approach, PECAN, to be significantly easier ($p<0.01$) and intuitive ($p<0.01$) than the indirect approach, APReL. This result supported our hypothesis H5.

### v2

@@ 6.1. Learning User-Friendly Canonical Spaces
Experimental Setup. Participants in this study interacted with a $6$-DoF UR5 robot arm in three manipulation tasks: (i) handing over a plate to the user (Handover Plate), (ii) placing a cup in front of the user (Place Cup), and (iii) pouring coffee into the cup (Pour Coffee). The states and actions for each task were the robot arm’s joint positions and velocities, respectively.
@@
The robot’s style was defined by the distance it maintained from the user. At one extreme, the robot could follow a straight-line path that comes very close to the user, while at the other extreme, it could take a curved path that stays as far from the user as possible. To personalize the robot’s trajectory, each user interacted with an interface containing a canonical space of the robot’s styles. The interface featured a task selection menu, a slider for choosing the latent style (as shown in Fig. 4-Left), and buttons to execute the trajectory corresponding to the chosen style in a Pybullet simulation and on the real robot. We pre-programmed the robot to pick up the objects for each task (i.e., plate, cup, or kettle) from their initial positions. Users first chose a task from the menu, which mapped to a latent task vector, then moved the slider to select the latent style. The robot then performed the task according to the style selected by the user. See videos here: https://youtu.be/wRJpyr23PKI
@@
Participants performed the three tasks in the order given above. In the first two tasks, Handover Plate and Place Cup, they could preview the robot’s trajectory in the Pybullet simulation before executing it in the real world. However, they did not have this option in the third task of Pour Coffee and had to find their target style based on their experience of using the interface in the first two tasks. We did this to determine if users could build an accurate understanding of the canonical space, enabling them to effectively transfer it to new tasks. Therefore, we treated the first two tasks as familiarization tasks and evaluated the user performances in the Pour Coffee task. We chose the Pour Coffee task for user evaluation because it was the most visually distinct among the three tasks, making it challenging for users to find their target style without visualization. We informed users that they would not be allowed to visualize the robot’s trajectory in simulation for the Pour Coffee task and must use the first two tasks to learn how the latent values mapped to the robot’s style. We anticipated that a consistent and monotonic canonical space would be easier for users to learn and apply across new tasks without the need to visualize the robot’s behavior.
@@
Independent Variables. We compared our proposed approach (PECAN) to a state-of-the-art baseline for generating latent style representations (SeGMA). To create the training dataset, we first provided three demonstrations each in the Handover Plate and Place Cup tasks. Two of the demonstrations in each task represented the extreme styles — trajectories with the maximum and minimum distance to the user. These $6$ demonstrations were common across all participants. In addition to the demonstrations provided by experimenters, we asked each participant to provide $2$ demonstrations in the Pour Coffee task, one for each of the extreme styles. In total, we trained custom PECAN and SeGMA models for each participant using a dataset of 8 demonstrations along with labels for the trajectories representing the extreme styles. For PECAN, we asked participants to assign the same label to trajectories with similar styles across the tasks. On the other hand, an expert provided the same label to trajectories in the same task for SeGMA. We modeled both approaches to have a $d_{\theta}=1$ dimensional latent style space and $d_{\tau}=3$ latent tasks.
@@
Table 1. Survey questions (Likert scales with $7$-option response format). We grouped the questions for the robot study into five scales: Easy, Intuitive, Accurate, No Visuals, and Prefer. We included additional questions for the Learn scale in the autonomous car study. We tested the reliability of each scale using Cronbach’s $\alpha$. The reliability scores presented for the first five scales are based on the responses recorded in the robot study. The reliability score for the Learn scale is based on user responses in the autonomous car study.
@@
Learn
- I needed fewer tries to personalize the robot as I gained more experience with the interface. 0.81
- I needed more tries to personalize the robot as I gained more experience with the interface.
Participants and Procedure. We recruited $14$ participants ($3$ female, average age $28\pm 5$ years) from the Virginia Tech community. Participants gave informed written consent under IRB #23-1237. At the start of the experiment, we informed the participants that the robot’s style refers to the distance it maintains from their body and asked them to kinesthetically guide the robot arm to demonstrate trajectories for the extreme styles in the Pour Coffee task. We then trained both methods on the user demonstrations along with the six previously collected demonstrations. Users interacted with the robot in all three tasks — once with each method. They completed the tasks in a fixed order but the ordering of the methods was counterbalanced: half of the users started with PECAN and the other half started with SeGMA.
In the Handover Plate and Place Cup tasks, users personalized the robot’s trajectory once to match distinct target styles. Then, in the Pour Coffee task, users personalized the robot’s motion twice to achieve two additional target styles. We randomly sampled the four target styles for each user. These styles were different from the demonstrations in the training data and could correspond to any point in the canonical space. For each style, users had $3$ attempts to perform the task on the real robot while ensuring that its trajectory stayed within a small tolerance of the desired distance. To help users gauge the actual style of the robot’s trajectory, we displayed its maximum distance from the user on the interface.
Dependent Variables. For each task we recorded the number of attempts that users took to achieve the target style on the real robot (Real Attempts). We also recorded the number of times users visualized the latent styles in simulation before executing them in the real world (Sim Attempts) for the Handover Plate and Place Cup tasks. A higher number of attempts indicated that it was difficult for users to identify their desired styles in the learned canonical space. Specifically, in the Pour Coffee task, where users did not have the option to simulate the styles, a higher number of attempts indicated that the interface was not intuitive and consistent across tasks. We also measured the absolute error in the distance (Style Error) and final position (Task Error) of the trajectories executed by users in their last attempt for each target style.
@@
Figure 5. Objective and subjective results from the robot user study. (Left) User applies the same latent value $z_{\theta}$ from Place Cup to Pour Coffee expecting a similar style across both tasks. PECAN produces trajectories with similar distances $\theta$ for Place Cup and Pour Coffee, but SeGMA generates a straight line trajectory for Pour Coffee. (Top Right) When using PECAN participants had lower task error, style error, and fewer real attempts (t(13) = -3.12, p < 0.01) performing new tasks without visual information. (Bottom Right) Subjectively participants prefer (t(13) = 2.55, p < 0.05) working with PECAN compared to SeGMA, as they found PECAN more intuitive (t(13) = 3.35, p < 0.01), and easy (t(13) = 2.68, p < 0.05) especially without visual information (t(13) = 2.76, p < 0.05). An asterisk (*) denotes statistical significance and the error bars indicate standard error.
@@
In the second user study we determined the pros and cons of directly selecting the robot’s style compared to indirect methods that estimate the style from user feedback (e.g., ranking robot trajectories). Here we shifted to the driving environment which is a standard benchmark for inferring user preferences and enables us to test the personalization of more complex 2D styles, as opposed to the simpler 1D style in the robot manipulation task. Participants were presented with two fundamentally different approaches for modifying an autonomous car’s driving style — our direct approach, PECAN, and an established approach for indirectly learning from human preferences (Sadigh et al., 2017).
Independent Variables. Specifically, we compared PECAN to an active preference-based learning approach, which we refer to as APReL. We implemented this approach based on the code provided in (Bıyık et al., 2022b). At each step, APReL presented two trajectories to the users, each representing a different style, and asked them to choose their preferred trajectory. Based on their choices, APReL inferred the individual user styles. It strategically selected these trajectories to maximize the information it gained about the user’s style from their choice. For example, showing trajectories with distinctly different speeds, such as one high-speed and one low-speed, is more informative than showing two trajectories with similar speeds. In our implementation, APReL selected the trajectories from a discrete space of $176$ uniformly sampled styles. We trained PECAN with $24$ demonstrations: $8$ demonstrations corresponded to the extreme styles and $16$ were sampled randomly, as explained in Section 5. We used a $d_{\theta}=2$ dimensional latent style space for PECAN and, correspondingly, $2$ dimensional features for APReL. It is important to note that, unlike our approach, APReL has direct access to the features that parameterize the driving style of the autonomous car. Therefore, we might expect this baseline to outperform our approach because it knows the actual styles, while our approach must learn these styles from user demonstrations.
@@
Dependent Variables. For the indirect approach (APReL) we recorded the total number of queries that users had to answer for personalizing the car’s behavior. For the direct approach (PECAN) we counted the total number of points that users had to visualize for finding their preferred style. We refer to the total number of queries or clicks required to achieve the target style as Attempts. Although clicks and queries represent different actions, they indicate the number of times users had to make a decision: where to click or which trajectory to choose. We also measured the difference (i.e., Euclidean distance) between the target style and the style achieved using each approach (Style Error). We did not compare the total time required to achieve the target style due to the fundamental differences between the two approaches. While PECAN is trained offline and instantly decodes the selected style, APReL requires significantly more time as it generates the queries and learns the styles online.
Finally, to subjectively compare the direct and indirect approaches, we collected user responses for the same scales as in Table 1. We also added another scale for measuring if users perceived that they required fewer attempts (clicks or queries) to personalize the robot as they gained more experience with each method (Learn).
@@
Figure 6. Objective and subjective results for the second user study. (Left) The average error in the style of the personalized car trajectory in both tasks. (Center) The total number of attempts (queries or clicks) required by users to personalize the car in both tasks. (Right) Users found both the direct and indirect approaches to be intuitive, accurate, and easy to use. In particular, they perceived that they needed fewer clicks to find their desired style as they gained more experience with our direct approach. An asterisk (*) denotes statistical significance and the error bars indicate standard error.
@@
Figure 7. Objective and subjective results for the follow-up study comparing direct and indirect approaches for personalizing robot styles with experienced users. (Left) The average error in the style of the personalized car trajectory in both tasks. (Center) The average number of attempts (queries or clicks) required by users to personalize the car for each target style across both tasks. After practicing with both approaches for $10$ minutes, users were able to achieve their target style more accurately with our direct approach, PECAN, and needed fewer attempts to do so than the indirect baseline, APReL. (Right) Subjectively, experienced users found PECAN to be more intuitive, accurate, and easy to use than APReL. Consequently, they preferred it over the indirect baseline for personalizing the car’s trajectory across tasks. An asterisk (*) denotes statistical significance and the error bars indicate standard error.
@@
Practice. Many participants demonstrated a similar approach to finding their preferred style using PECAN. They first visualized the car’s trajectory for two to four points in different regions of the canonical space to understand how the latent dimensions aligned with the actual styles. Based on this initial exploration, participants chose a point that would most probably result in their desired style. If this point was still outside the allowed limits, they incrementally moved the point in a direction that would bring them closer to their target. We did not observe such guided behavior when users practiced with APReL. To assist users in recalling their initial interactions, we displayed their previous clicks in PECAN’s canonical space and their previous query responses on APReL’s interface. However, some users still preferred to note down the actual styles corresponding to their clicks or query responses on paper, suggesting that both interfaces could benefit from more intuitive visual cues.
Results. The results of the follow-up study are shown in Figure 7. After practicing with both approaches, we observed that users required significantly fewer clicks or queries to personalize the autonomous car’s style with PECAN than with APReL. When using PECAN, many participants were able to skip the initial exploration phase that they exhibited during practice. A two-tailed paired t-test showed a statistically significant difference ($t(9)=-3.28$, $p<0.01$) between the average number of attempts required by users per style with PECAN ($M=2.1$, $SE=0.33$) and APReL ($M=3.9$, $SE=0.48$). Subjectively, experienced users found our direct approach, PECAN, to be significantly easier ($p<0.01$) and intuitive ($p<0.01$) than the indirect approach, APReL. This result supported our hypothesis H5.
@@
Applications. Both APReL and PECAN offer distinct advantages. PECAN enables users to instantly choose their desired style, but requires them to determine the specific point that aligns with their preference. While it places some cognitive burden on users, our follow-up study indicates that this burden decreases with practice. In contrast, APReL shifts the cognitive burden to the robot, which selects the trajectories to present to the user and infers the user’s preference based on their choices. This reduces the user’s immediate effort but requires significantly more time and interactions for the robot to learn preferences online. Another distinction is that PECAN learns a fixed canonical space, while recent approaches have extended APReL to incorporate new features through novel forms of human demonstrations (Bobu et al., 2022). In our current approach, users would need to demonstrate and label the extremes of any new style variable they wish to add to PECAN’s canonical space.
Overall, we posit that PECAN is best suited for applications where users want to quickly and repeatedly change the robot’s style within a fixed spectrum. For example, users may want to incrementally make their autonomous car more aggressive or defensive depending on how late they are running. Conversely, since APReL cannot facilitate quick adjustments, it may be preferable in scenarios where users want to passively respond to the robot’s queries, aiming to define the robot’s style once and then repeatedly use the same style.
7. Practical Considerations
In this section, we discuss the key design considerations for implementing and using PECAN in practice. Here we also mention PECAN’s limitations and potential directions for future work.
Dimensionality of canonical space. In our experiments, we assume that the dimensionality of the styles is known and design the canonical space to have the same number of dimensions as the true styles. But more generally, the dimensionality of the canonical space does not need to match the true style variables. Rather, it depends on the number of style variables that the user wants to personalize. For example, the aggressiveness of an autonomous car can depend on multiple variables such as the car’s speed, the distance it maintains from other vehicles, the number of times it changes lanes to overtake other cars, etc. However, users may not want to tune each variable individually — they may simply care about changing the overall driving style to be more aggressive or defensive. In this case, we would only need a 1D canonical space with the most aggressive driving style at one end and the least aggressive style at the other.
Therefore, practitioners can decide the dimensionality of their canonical space based on the number of extreme styles labeled by the user. Ideally, users must label $2^{d}$ extremes for a $d$-dimensional latent space, e.g., a 3D style space would have $8$ extremes (i.e., corners of a cube). Using a canonical space that is smaller or larger than the exact dimensions needed to model the extreme labels can significantly affect PECAN’s performance. Our experiments in Appendix A.2 demonstrate that adding an extra dimension causes the latent styles to also encode task information, decreasing the consistency of the canonical space. Conversely, removing a dimension of the canonical space preserves consistency, but it prevents the styles from being monotonically arranged in the canonical space. Future work can explore data-driven approaches that do not rely on having access to extreme styles for determining the dimensionality of the canonical space (Kobayashi and Watanuki, 2023).
Demonstration of extreme styles. A key feature of PECAN is the monotonicity of its canonical space which makes the interface intuitive to use. By minimizing the combined loss in Equation 6, PECAN encodes the extreme styles to the corners of the canonical space, while placing the intermediate styles monotonically between the extremes. To achieve this, PECAN must have access to the demonstrations and labels of all the extremes of the style spectrum. When demonstrations and labels for certain extremes are missing, the cross-entropy loss will still form a simplex with the available labels. For example, if the data only contains labels for two of the four extremes of a 2D style space, PECAN will encode them to opposite corners of the canonical space leaving the other two corners to potentially encode intermediate styles. Therefore, the learned canonical space may no longer be monotonic.
Our experiment in Appendix A.2 demonstrates that the monotonicity of the canonical space reduces when the extreme labels do not align with the dimensionality of the canonical space. To address such cases with our current implementation, practitioners would either need to reduce the dimensionality of the canonical space to match the existing labels or provide the missing labels for extremes of the available demonstrations. We emphasize that labeling the style extremes is more intuitive than previous approaches which require users to quantify the robot’s style (Schrum et al., 2024). Developing unsupervised methods that can automatically identify the style extremes by comparing the trajectories is an exciting direction for future work.
Reconstructing intermediate styles. In addition to demonstrations of extreme styles, PECAN also requires unlabeled demonstrations of intermediate styles for training the decoder to accurately reconstruct trajectories from different latent values. Our preliminary tests showed that when the data does not contain a sufficient number of intermediate demonstrations, the decoder may fail to generate meaningful trajectories for certain points in the canonical space. In our experiments, we mitigated this concern by providing the minimum number of demonstrations required to train the decoder effectively. We found that PECAN requires more training demonstrations as the complexity of tasks and styles increases. For instance, $8$ demonstrations were sufficient to learn the manipulation tasks in our user study because the environment was static and the styles were only one dimensional. On the other hand, we needed $16$ demonstrations to learn the dynamic driving tasks which had two-dimensional styles. We have conducted further experiments in Appendix A.3 demonstrating that PECAN’s reconstructions improve as we increase the number of intermediate styles in the training data. Future work should investigate how the decoder’s output can be constrained to guarantee that the generated trajectories are safe even when trained with insufficient demonstrations.
Personalizing high-dimensional styles. In this work, we evaluate PECAN across styles ranging from one to three dimensions. As explained earlier, the dimensionality of the latent styles depends on the number of variables that users choose to personalize, rather than the total number of variables that define the robot’s behavior. Typically, we expect that users would only prefer personalizing a few style variables directly. However, can PECAN retain its user-friendly properties when learning canonical spaces with more than three style dimensions? To explore this, we present additional experiments in Appendix A.4 testing PECAN on styles with four to six dimensions. Our results indicate that PECAN maintains its monotonicity as we increase the dimensionality of the canonical space with only a small decrease in the consistency of its canonical spaces.
Another consideration when learning high-dimensional styles is how the canonical spaces can be visualized for direct personalization. In our experiments, we present two interfaces for visualizing the canonical space: a linear slider for 1D styles and an interactive square for 2D styles, as shown in Figure 4. Practitioners can use a combination of these interfaces to represent canonical spaces with three or more dimensions. For example, we can use $n$ sliders to personalize $n$-dimensional robot styles. In such cases, we may want each latent dimension (i.e., slider) to correspond to a distinct style variable so that the interface is intuitive to use.
Prior research in representation learning has explored the problem of disentangling the latent variables such that each dimension represents an independent factor of variation (Wang et al., 2024). Many of these approaches use Variational Autoencoders (VAEs) to disentangle the latent dimensions in an unsupervised manner by forcing the posterior latent representation to be a multivariate Gaussian with an Identity covariance matrix (Burgess et al., 2018; Mathieu et al., 2019). However, prior studies have shown that while this disentangles the dimensions of the posterior samples, the dimensions of its mean (which is what users will specify when choosing their desired style) can remain correlated (Locatello et al., 2019). Hence, some form of supervision is necessary to disentangle the latent dimensions into independent variables that the user wants to personalize (Hristov et al., 2020).
While our current approach obtains labels for the extremes of the style space, it does not know which extremes correspond to opposite ends of the same style dimension. As a result, the dimensions of our canonical space can get entangled. Therefore, when learning high-dimensional canonical spaces, practitioners can obtain labels for extremes of each style dimension individually. Although this requires humans to have additional domain knowledge, they would only need to label $2\times d$ extremes for a $d$-dimensional style, which is significantly fewer than the $2^{d}$ labels required in our current implementation. For example, for a 3D style $\theta$ = [$\theta_{1},\theta_{2},\theta_{3}$] users will need to label $6$ extreme styles, i.e., the low and high values for each dimension — [low $\theta_{1}$, high $\theta_{1}$], [low $\theta_{2}$, high $\theta_{2}$], and [low $\theta_{3}$, high $\theta_{3}$]. In contrast, as discussed earlier, our current approach would require $8$ labels (i.e., corners of a cubic latent space). With these labels, we can train separate 1D canonical spaces for each style variable using the same loss functions proposed in Section 4. We look forward to investigating such approaches for disentangling the latent dimensions in future work and examining whether users would find it intuitive to personalize more than three style dimensions.
Other variations in trajectories. Lastly, PECAN assumes that all tasks in the dataset share the same styles. In our simulation experiments, we found that as the tasks become more dissimilar, PECAN finds it easier to distinguish their trajectories and encode the latent task vectors. However, in real-world scenarios, when tasks become too dissimilar they may no longer have the same styles. In such cases, we may need to identify subsets of tasks that share similar styles and then learn separate canonical spaces for these subsets.
Another limitation of our approach is that we only consider variation in trajectories due to styles specified by the user. However, in many tasks, trajectories can vary due to environmental factors such as changes in object positions or the behavior of other agents. To account for such variations, we would need to condition our proposed architecture on the environment state. For example, we can provide the object positions as an additional input to the style encoder and trajectory decoder to separate the variation in styles from the environment variations in a robot manipulation task. In this work, we assume that the environment remains consistent to emphasize the personalization of robot styles and leave extending our approach to include environment variations for future work.